Solve system by the substitution method and indicate wheter each system is independent, inconsistent or dependent
3y=x+5
3x-9y=-10

Answer by stanbon(75887)

Solve system by the substitution method and indicate wheter each system is independent, inconsistent or dependent
3y=x+5
3x-9y=-10
------------
Multiply thru the 1st equation by 3 to get:
9y = 3x+15
Substitute that into the 2nd equation to solve for "x":
3x -(3x+15) = -10
-15 = -10
-----
The system is inconsistent; therefore, it has no solutions.
=================================================
Cheers,
Stan H.